OpenStudy (anonymous):

3x^2-5x-12=0 solve the equation check your solution show work

OpenStudy (anonymous):

3x^2 -9x + 4x -12 = 0 3x(x-3) +4(x-3)=0 (3x+4)(x-3)=0 hence 3x+4 = 0 so x = -4/3 and x-3 = 0 so x =3 so x = -4/3, 3

OpenStudy (anonymous):

3x²-5x-12=0 3x² - 9x + 4x - 12 = 0 3x (x - 3) + 4 (x - 3) = 0 (3x + 4) (x - 3) = 0 x = 3 or x = -4/3

OpenStudy (amistre64):

(x- /3)(x+ /3) 1,36; 2,18; 4,9.... (x-9/3)(x+4/3) (x-3)(3x+4) x = 3; x=-4/3 correct :)